OpenStudy (gabylovesyou):

Based on the graph of y = x2 − 2x − 3, what is the x coordinate of the positive x-intercept?

OpenStudy (gabylovesyou):

@dumbcow

OpenStudy (aravindg):

The x intercepts will be the roots of the given quadratic

OpenStudy (mertsj):

Find the x intercepts by replacing y with 0

OpenStudy (aravindg):

Out of the 2 roots .The one which is positive will be the x intercept you are looking for.

OpenStudy (aravindg):

Got it?

OpenStudy (gabylovesyou):

0 = x^2 - 2x - 3 what do i do after this @Mertsj

OpenStudy (gabylovesyou):

@AravindG how do i solve it ?

OpenStudy (mertsj):

I think you are supposed to graph it

OpenStudy (aravindg):

yes^^

OpenStudy (mertsj):

The problem says...based on the graph

OpenStudy (mertsj):

So do you know how to find the vertex?

OpenStudy (aravindg):

Or you just factorize :)

OpenStudy (gabylovesyou):

vertex is x = -b/2a right ?

OpenStudy (mertsj):

yes. That is the x coordinate of the vertex.

OpenStudy (gabylovesyou):

ok i will solve it now hold on

OpenStudy (mertsj):

ok

OpenStudy (gabylovesyou):

is it x = -(-2) / 2(1) ??

OpenStudy (mertsj):

Yes. so x = 1

OpenStudy (gabylovesyou):

now ill solve for y hold on

OpenStudy (gabylovesyou):

y = -4

OpenStudy (mertsj):

Graph that point. (1,-4)

OpenStudy (gabylovesyou):

now what

OpenStudy (mertsj):

Replace x with 0 and find y

OpenStudy (gabylovesyou):

but to what equation ? lol

OpenStudy (mertsj):

The only equation we have. y=x^2-2x-3

OpenStudy (mertsj):



OpenStudy (gabylovesyou):

y = -3

OpenStudy (gabylovesyou):

right @Mertsj

OpenStudy (mertsj):

Then use symmetry to find the corresponding point to the right of the vertex...as shown in the drawing.

OpenStudy (mertsj):

Now let x = -1 and find two more points.

OpenStudy (gabylovesyou):

y = 0

OpenStudy (mertsj):